OpenStudy (anonymous):

simplify. 1/(x+h)^2-1/x^2/h solved . just want to check if i did right lol

OpenStudy (anonymous):

Show your work first...

OpenStudy (anonymous):

[1/(x+h)^2-1/x^2/h=1/(x+h)^2/h/-1/x^2/h\] \[1/h(x+h)^2-1/x^2h\] \[x^2/x^2h(x+h)^2-(x+h)^2/x^2h(x+h)^2=1/h\] that;s what i got.... =_=;;

OpenStudy (anonymous):

oh -1/h